solve the following system using substitution
y=-3x+4
y=2x-11

Answer by ankor@dixie-net.com(22740)

y=-3x+4
y=2x-11
:
Replace y in the first equation by 2x-11
2x - 11 =-3x + 4
2x + 3x = 4 + 11
5x = 15
x = 3
:
You can find y